OpenStudy (anonymous):

Help simplifying the expression -5 + i/2i

OpenStudy (freckles):

is this (-5+i)/(2i)?

OpenStudy (freckles):

\[\frac{-5+i}{2i}?\]

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

ok and the objective i bet is to put into a+bi form

OpenStudy (anonymous):

I think so

OpenStudy (freckles):

we know that i*i=i^2=-1 so to get rid of that imaginary unit in the denominator multiply both top and bottom by i

OpenStudy (freckles):

\[\frac{-5+i}{2i} \cdot \frac{i}{i} =?\]

OpenStudy (freckles):

what is (2i)*i=? and what is (-5+i)*i=?

OpenStudy (anonymous):

-5i -1 / -2 ?

OpenStudy (freckles):

\[\frac{-5i+i^2}{2i^2}=\frac{-5i-1}{-2}=\frac{-1-5i}{-2}\] now there is a couple of more things to do to make it prettier

OpenStudy (freckles):

like all the terms on top and all the terms on bottom all have a -1 factor in common

OpenStudy (freckles):

\[\frac{(-1)(1+5i)}{(-1)(2)}=\frac{1+5i}{2}=\frac{1}{2}+\frac{5}{2}i\]

OpenStudy (freckles):

good job

OpenStudy (anonymous):

i think im suppose to give and answer

OpenStudy (freckles):

what do you mean

OpenStudy (anonymous):

the actual answer to -5 + i/2i not just putting it into a+bi form

OpenStudy (freckles):

a+bi is usually the standard form to put it into

OpenStudy (anonymous):

The answer was -4.5 how do I get that

OpenStudy (freckles):

so when I asked you if it was \[\frac{-5+i}{2i}? \] you didn't mean to answer yet right?

OpenStudy (freckles):

and you really meant \[-5+\frac{i}{2i}?\] there is a difference between these fractions

OpenStudy (anonymous):

oh okay nvm

OpenStudy (freckles):

\[-5+\frac{i}{2i} \\ -5+\frac{i}{2i} \cdot \frac{i}{i} \\ -5+\frac{i^2}{2i^2} \\ -5+\frac{-1}{2(-1)}\] I think you can finish simplifying this

OpenStudy (freckles):

you could have just canceled the i's in the first step -5+(1/2)